OpenStudy (anonymous):

give the zero's of this problem pls. F(x) = (2x+1)(x-5)^2

OpenStudy (nenadmatematika):

x=-1/2 x=5

OpenStudy (anonymous):

thats all of the zeroes ? :) and thanks

OpenStudy (nenadmatematika):

well yes, but you can say if your professor asks you to, that that zero x=5 is double so you would write in that case that the zeros are x=-1/2 x=5 and x=5

OpenStudy (anonymous):

thank u buddy. and i aint talkin to a professor its for my hard retricemath 3 class online :p but i got another question if ur available

OpenStudy (nenadmatematika):

OK :D

OpenStudy (anonymous):

f(x)=(x-1)^3+2 describe the transformation